Shift summary for evening shift 5 Sept 99

Sasha Glazov (glazov@hep.uchicago.edu)
Mon, 6 Sep 1999 00:25:45 -0500

Shift Summary for eve shift, 5/Sept/99

On Shift: VO, SG.

Hours of beam available at all intensities: 7:40min
Hours of beam available at nominal intensity: 7:40min
Hours of E832 data (special runs not included): ~7:20min
Hours of special runs (laser scans, muon runs...): laser scan

Accounting of lost time: Tape change, kfrends

========================================================================
Runs ended this shift:
Run 14431 -- 23.0M beam_e832
Run 14432 -- 21.6M beam_e832
Run 14433 -- Laser_scan (1 pass)

Tapes brought to FCC this shift: QKx393-395, QKB138
(including tapes for online split)
none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking. At 18:32 HA2 scaler rate jumps twice. After
investigations with ET,EB & RK we believe that it is due to bad scaler -->
no problem for data taking. Scaler to be fixed next down time.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14431 16:00 Tapes are full, run ends
14432 16:08 Start new run (after 1 kfrend)
16:26 SEM is >= 7.0E12 for ~10 spills causing DAQDEAD
alarms, call MCR --> helps
18:10 DN: 2 new bottles of argone ethane on line east side
105 18:32 Alarm HA2/RegTop = 0.23 (limit 0.15) goes off.
Related to HA2 scaler rate which becomes twice
higher than normal. Paged Elizabeth, Ed, Rick.
Elizabeth comes. After many investigations
it seems like the problem is due to a bad scaler
module for HA2 --> all other things look normal.
See log book for details. VO changes alarm
limits to 0.25.
145 19:20 Ktevana errors --> 1 kfrend
194-196 20:24 ktevana errors, ktrig window does not respond,
restart, kfrend --> Ok.
227-229 21:10 live time=0 alarm. Busy in banana crates, reset
kumquats, kfrend --> ktevana errors, one more
kfrend --> Ok.
260-- Intensity is down to ~6E12. Called MCR. They say
they are tuning.
285 22:35 Intensity is back to normal
23:12 TV monitor message: Quench at A1 beam estimate
2 hours. Stop the run to make laser scan.
Paged Ed if we should attempt fix of HA2 scaler.
He suggest to leave it as it is for now.
VO is fixing her car.
14433 23:25 Start laser scan. Oops, 1 pass only
14434 23:50 Start looong laser scan.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Do not do muon run.